Let n be a natural number. Use mathematical induction to prove that
4
n−1 > n2
for all n ≥ 3.
1
Expert's answer
2020-06-10T19:51:00-0400

Let the statement "P(n) = 4^{n\u22121} > n^2"

Now, "4^{3-1} = 4^2 = 16 > 3^2" , So "P(n)" is true for "n = 3" .

Let "P(n)" is true for "n = k" "\implies 4^{k-1} > k^2" .

Now, prove "P(n)" is true for "n = k+1."

"4^{(k+1)-1} = 4 \times 4^{k-1} > 4 \times k^2 > k^2 + 2k^2+k^2 > k^2 + 2k +1" because "k \geq 3" .

Hence, "4^{(k+1)-1} > (k+1)^2"

Hence, using Principal of Mathematical Induction, "4^{n-1}>n^2" for all "n\geq 3" .